OpenStudy (anonymous):

can you explain me to me how to graph slopes? example:slope 6 intercept 20 y=6x+20

OpenStudy (anonymous):

*

OpenStudy (anonymous):

The slope describes the rise over the run type of path that each point follows, and the 20 describes the y intercept

OpenStudy (anonymous):

Probably one of the simplest methods to draw the graph would be to find the x and the y intercepts

OpenStudy (anonymous):

Are you familiar with being able to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I know that part but the I don't get the part where I move over and go down or up I forgot what to do.

OpenStudy (anonymous):

As long as you are able to get two points from the equation, then you can draw the graph, this being one way to draw a graph when given a formula like this

OpenStudy (anonymous):

What you are talking about is the slope when you have a point on the graph. The slope is the 6 and describes the rise over the run of how the graph moves

OpenStudy (anonymous):

Rise being the increase or decrease in the y direction and the run being the increase or decrease in the x direction

OpenStudy (anonymous):

ohh okay is there suppose to be 1 under the 6 since its the slope

OpenStudy (anonymous):

Yes exactly

OpenStudy (loser66):

@jashi34 we don't graph the slope, we graph the line, and that line "slant" the amount of slope All you have to do is jot down the equation of the line, then, let x =0 to have y, and let y =0 to have x. Link 2 points, you get the graph with the slope you want

OpenStudy (anonymous):

so I just graph 6 and 20

OpenStudy (amistre64):

y=6x+20 ; when x=0, y=20, the point (0,20) is plotted you say your confused how to use the slope to find another point. think of the slope as a fraction: 6/1 which relates to (the change in y) / (the change in x) the change in y .... y changes by 6 giving us: 20+6 = 26 the change in x .... x changes by 1 giving us: 0+1 = 1 therefore, another point on the line is defined as (1,26) we can find many points on the line if we just follow the slope, they can be defined as: the change in y .... y changes by 6n giving us: 20+6n the change in x .... x changes by 1n giving us: 0+n notice that this gives us the equation of y back to us: y = 6n + 20 for some multiple of n is the same statement as y=6x+20 for some value of x. later on you will find these to be parametric equation: x = n y = 6n + 20